You will be using the function f(x) = 1/x

1) Describe the shape of a typical cross-sectional slice of the solid and then write down the formula of the area of that cross section.
2) Use an integral to express the total volume of this solid between x=1 and infinity. Explain why the integral represents the volume described.
Answer #1

lets take a randon solid whose cross is eircllar. And surface of solid is rotalingy.= f(x) aberit x-azis section generated arf() lets take x70 x-axis, la surface @ when we rotate flas about is generated.c the slice (cross sections / will be a circulaV Eseblaination ich we rotate y = f(x) about & aris we get crow section as circle area of circle = 72 = a(fa) 2 then if we ke